# A unified framework for coarse grained molecular dynamics of proteins

### v1

@@ Abstract
Understanding protein dynamics is crucial for elucidating their biological functions. While all-atom molecular dynamics (MD) simulations provide detailed information, coarse-grained (CG) MD simulations capture the essential collective motions of proteins at significantly lower computational cost. In this article, we present a unified framework for coarse-grained molecular dynamics simulation of proteins. Our approach utilizes a tree-structured representation of collective variables, enabling reconstruction of protein Cartesian coordinates with high fidelity. The force field is constructed using a deep neural network trained on trajectories generated from conventional all-atom MD simulations. We demonstrate the framework’s effectiveness using the 168-amino protein target T1027 from CASP14. Statistical distributions of the collective variables and time series of root mean square deviation (RMSD) obtained from our coarse-grained simulations closely resemble those from all-atom MD simulations. This method is not only useful for studying the movements of complex proteins, but also has the potential to be adapted for simulating other biomolecules like DNA, RNA, and even electrolytes in batteries.
@@ I Introduction
Within the bottom-up paradigm of CG methods, researchers leverage the framework of statistical mechanics to define effective interactions, such as potential energiesLiwo et al. (2001). Over the past decades, several traditional coarse-grained methods have emerged, demonstrating significant improvements in capturing protein dynamicsMim et al. (2012); Chu and Voth (2005); Yu et al. (2021); Tóth (2007); Li and Takada (2010). The incorporation of machine learning (ML) techniques has further enhanced the performance of CG simulationsYu et al. (2021); Wang et al. (2019); Husic et al. (2020). For instance, Zhang et al. employed deep neural networks (DNNs) to approximate potential energy surfaces from coarse-grained coordinates, achieving consistent distributions of collective variables compared to those obtained from original MD simulationsZhang, Wang, and Weinan (2018). This framework further enabled enhanced sampling of large proteins Wang et al. (2022).
@@
In this research, we present a unified and recursive framework for constructing a bidirectional map between protein Cartesian coordinates and a minimal set of interpretable coarse-grained collective variables. This framework incorporates all heavy atoms in the backbone and side chains, enabling a comprehensive representation of the protein structure. Notably, our approach allows MD simulations to be performed solely using the collective variables, eliminating the need for explicit Cartesian coordinates. This significantly accelerates simulations while maintaining consistent performance with all-atom MD simulations, as measured by collective variable distributions and RMSDs at most time steps. Compared to existing methods that achieve only overall statistical consistency on collective variablesZhang, Wang, and Weinan (2018), our framework offers the advantage extra high accuracy in time series of RMSDs.
@@ II.1 Collective variables
Our protein representation begins with a general description of atomic coordinate hierarchy in chemical compounds. For clarity, this work adopts a non-standard notation. The symbol will represent any dihedral angle and will represent any bond angles. According to referencesParsons et al. (2005), the operating matrix for rotating along a normalized rotation vector is
where is the identity matrix, and . Considering the transformation, the general formula of coordinate transformation goes as a matrix, where is a vector filled with zeros and is a matrix representing the translation.
Converting the coordinates in local axis frame (x’y’z’) to parent axis (xyz) can be understood as transforming the parent axis data (axis vector and the origin) to the the local axis data in the parent axis frame, as illustrated in Fig. 1.
@@
, where first, translates the frame along x axis by bond length where and ; second, rotates the frame along the new z axis by bond angle with , and ; third, rotates the frame along the new x axis by dihedral angle with , and .
@@
with index showing the coordinate frames before current frame, and and being the target and current frame index. We denote the global operator for converting current coordinates to the "global" reference as
@@
can be readily applied to determine any protein atom’s Cartesian coordinates within the global reference frame, where vector and are original Cartesian coordinates () and extended coordinates (). For convenience, subscript indicates the number and index of coordinate transformation from the "global" coordinate. A superscript of "0" denotes a "local" coordinate within its frame, otherwise it’s a "global" coordinate. As are only determined by the amino type, the Cartesian coordinates are only dependent on . The bond lengths, , used in the calculations can also be retrieved from a the standard amino acid structures. It is noteworthy that the connections between amino acids in a protein’s backbone, specifically between the nitrogen and carbon atoms, have a fixed length of about 1.32 angstroms.
The conversion between a protein’s 3D structure (Cartesian coordinates) and its internal torsion and bond angles is straightforward and is neglected here. This establishes a bidirectional mapping, facilitating analysis from both perspectives. To optimize computational efficiency, calculations consider only the essential heavy atoms. Additionally, rigid moieties like benzene rings require only two angles for complete description. This significantly reduces the number of variables needed, streamlining computational demands. While introducing some additional parameters compared to traditional methods, our approach offers substantial advantages. It allows reconstruction of the complete 3D structure, including side chains, leading to more accurate force field predictions – crucial for molecular simulations.
@@ II.1.2 Tree data structure
Recursive data structures naturally lend themselves to representation using tree structures. This approach has been successfully implemented in J. Zhu’s prior work on quantum physics Zhu (2021); Zhu and Scrinzi (2020); Zhu (2020). Similarly, the coordinates of each reference frame within a protein can be effectively represented as a tree structure, as demonstrated in Fig. 1. In this tree, each data node inherits its information from its parent node. This tree-based representation not only captures the hierarchical relationships between protein components but also partially incorporates the protein’s geometric information. For a more detailed breakdown of the hierarchy within each amino acid, please refer to Table 1 in the appendix.
@@
While the recursive multiplications of matrices can be computationally expensive, the reconstruction of protein 3D structures is only required during trajectory analysis. This means these computationally intensive operations are not necessary for real-time molecular dynamics simulations, thereby minimizing their impact on overall performance. Furthermore, the reconstruction of collective variable trajectories can be efficiently parallelized across CPUs and GPUs. This distributed processing approach effectively mitigates the performance drawbacks associated with recursive computation time.
II.2 Force field and molecular dynamics
Simulating large biomolecular systems necessitates capturing their statistical behavior. There exist two main approaches exist: stochastic methods and equilibrium sampling. Stochastic methods introduce random noise at each simulation step to mimic statistical fluctuations, starting from a defined initial state. Examples include Monte Carlo, random walk, and diffusion methods. Equilibrium sampling uses an initial state that already reflects the desired statistical properties, followed by noise-free simulations. This approach is employed in traditional MD simulations. Our work aligns with the second strategy (equilibrium sampling) by utilizing a machine learning force field trained on MD outputs. In contrast to methods that learn the energy landscape from existing trajectories, our approach directly predicts the next MD coordinate. This is possible because the statistical uncertainty is already embedded within the initial velocity distribution used during training. Subsequent steps involve solving Newton’s equations of motion, efficiently leveraging the learned velocity information. The network’s ability to predict the next coordinate while capturing statistical behavior becomes evident in the loss function design (Eq. 13, ). This design trains the network on a sequence of consecutive frames, effectively capturing velocity and force field information from training data. Consequently, the task becomes one of predicting the next CVs based on the previous step, as the network has already learned the underlying statistical dynamics.
@@ II.2 Force field and molecular dynamics
The network’s task is now to predict the subsequent collective variables based on the current ones. These predicted variables are also projected to sine-cosine values, formulated as
@@
, where denotes the number of hidden layers in the network, and represent the time indices (timestamps) of the trajectory data. The formula of function in each layer is written as
@@
, where and are weights to be learnt, and are activation functions. Our network distinguishes itself from Zhang et.al’s Zhang, Wang, and Weinan (2018) work by introducing a normalization constraint,, denoted by , for the sine-cosine operation within the network. This normalization step helps mitigate potential artifacts arising from the projection of angular data into sine and cosine components. While these constraints could also be incorporated into the loss function, we found that applying them directly through significantly accelerates the convergence of the network during training.
@@
The task is to obtain the weights that minimize the loss that can be described as . We use mean square error as the loss function which goes as
@@
, where we use error between variables at and those obtained from . This loss can also be extended to a more general form as
, where the error computed from subsequent coordinates after are also used. It is easy to see Eq. 13 collapses to Eq. 12 with . Interestingly, when in Eq. 13, the network effectively learns the underlying physical form of the force field. The velocity-verlet algorithm, a popular method in molecular simulations, allows for efficient calculation of forces acting on particles. This method uses the particles’ positions at three time steps (t-1, t, and t+1) as detailed in
The calculation works regardless of the particles’ initial velocities. In Eq. 14, is the piece-wise multiplication operator, represents a constant mass assigned to each simplified unit (coarse-grained coordinate), is the predefined time step, and the acceleration vector is the only unknown variable to be solved for. Thus including data from at least three successive trajectories enables the network to learn the force field information.
@@ III.1 Protein Structure reconstruction
Fig. 3 depicts the reconstructed initial structure of protein T1027. The reconstructed tertiary structure (cyan) exhibits a near-native match with the original data (gray), highlighting the accuracy of our approach. For comparison, the blue structure was built using all angles except for the bond angles, which were fixed at their theoretical value (). This comparison reveals a significant mismatch, including the absence of part of the alpha helix. This demonstrates the substantial impact of even small variations in bond angles on protein structure, emphasizing the importance of including them in our representation. The secondary structure representation also shows good agreement between the reconstructed structure and the original data, with only minor deviations observed at the ends of a few amino acid side chains. Specifically, analysis of CA atom distances reveals a maximum deviation of 0.26 angstrom, with an average value of 0.04 angstrom. The maximum difference for side-chain atoms is 0.6 angstrom, with a mean of 0.26 angstrom. It is noteworthy that we utilize local coordinates (denoted as - see Section II.1) for each amino acid, retrieved from a separate database independent of the T1027 structure. This approach can introduce slight error accumulation along the protein chain backbone, which can then propagate to the side chains (whose positions are highly dependent on the backbone). However, as will be shown in subsequent sections, the errors introduced by our coarse-grained representation remain consistently below 0.2 angstrom in practical applications, making them effectively negligible.
@@ III.2 MD trajectories prediction
We employed the GROMACSVan Der Spoel et al. (2005); Lindahl, Hess, and van der Spoel (2001); Berendsen, van der Spoel, and van Drunen (1995) software (version 2022.2) for the MD simulations, utilizing a timestep of 1 femtosecond (fs) and saving trajectories every 0.05 nanoseconds (ns). To enhance the efficiency of the simulations, we implemented an in-house developed enhanced sampling method, parallel continuous simulated tempering and a structure based model (PCST+SBM), within GROMACS. For a detailed description of PCST+SBM, please refer to references Zhang and Ma (2010); Zang et al. (2014, 2018); Ma et al. (2018). We start the MD simulations from the refinement target in CASP14. Two copies are simulated in PCST, whose temperature distribution is defined by and , respectively. In configuring the SBM method, we selected the starting target structure as the reference to promote initial-state fidelity during the MD simulations. This ensures the dynamics primarily focus on exploring configurations near the initial structure. Compared to the original SBM implementation, we introduce a targeted constraint strategy. We constrain only the "slow" residues, which are crucial for calculating the RMSDs. Residues within loops are left unconstrained to facilitate efficient exploration of the conformational space.
@@
For training the deep neural network, we employ a common data split strategy: 70% of the processed trajectories are used for training, 20% for validation, and the remaining 10% for testing. We configure the network architecture with three hidden layers () and utilize two weight matrices () of size and (details on weight matrix dimensions might be specific to your field).
@@
Our DNN network exhibits remarkable efficiency. It can generate a 25ns trajectory using only the initial structure as input and applying the network recursively to predict subsequent coordinates. This process takes merely 0.59 seconds on a standard NVIDIA GeForce GTX 1660 Ti (6G) GPU. In stark contrast, traditional MD simulations on a 256-CPU supercomputer require roughly 20 hours to generate a comparable trajectory. This translates to an impressive speedup of approximately 10,000 times achieved by the DNN network. Even when considering the training time (around 2 hours), the coarse-grained DNN network offers a significant speedup factor of 10 for the entire process. It’s important to note that training can be further optimized by leveraging parallelization techniques and better GPUs.
@@
Fig. 4 depicts the predicted performance of the MD simulation. The upper panel shows the angles (torsion angles and bond angles ). As expected, the angles (represented in the first half) exhibit a wider distribution range due to their inherent flexibility in the protein backbone. In contrast, the angles (, shown in the second half) primarily cluster around 1.23 radians () with a smaller variance ( 0.4 radians) during the simulation. This finding reinforces the earlier observation of bond angle fluctuations within MD simulations. Notably, a small population of angles around radians likely originates from the carbonyl oxygen atoms in the backbone. The comparison of the orange (black) and red (predicted) curves reveals a high degree of consistency between the predicted statistical behavior of the angles and the original values. Similar observations can be made when analyzing the sine and cosine components of these angles, as presented in Fig. 5 in the appendix.
@@
The RMSD values calculated from the predicted CVs closely match those obtained from the original Cartesian coordinates (Fig. 4, black vs. red lines). Our observations align well with previous structural analysis, where deviations were consistently less than 0.2 angstrom and the average difference using angles () was 0.0691 angstrom with a standard error of 0.0418 angstrom. Additionally, the predicted trajectories closely resemble the original data, particularly between 15-25 ns and 0-5 ns for n=2, as shown by the RMSD profiles. This overall match is also satisfied for n=5 with better agreement during 0-5 and around 12 ns, while worse agreement around 10 and 16 ns. A convergence study (Fig. 6, Appendix) explored additional n values (1, 3). As expected, n = 1 resulted in significant deviations, indicating insufficient information is learned. Increasing n to 2 eliminated these deviations, suggesting successful information capture. Interestingly, n = 3 led to a slight performance decline, while n = 5 offered minimal improvement while potentially causing slower training and overfitting. Therefore, n = 2 is optimal for convergence. The observed convergence behavior aligns with our prior analysis (Eq. 14), where appears sufficient to capture the essential force field information and generate meaningful data for MD simulations. The close agreement between our predicted and reference RMSD values suggests a more accurate representation of protein dynamics in our coarse-grained method. This is particularly noteworthy because RMSD is highly sensitive to the choice of CVs.

### v2

@@ Abstract
Understanding protein dynamics is crucial for elucidating their biological functions. While all-atom molecular dynamics (MD) simulations provide detailed information, coarse-grained (CG) MD simulations capture the essential collective motions of proteins at significantly lower computational cost. In this article, we present a unified framework for coarse-grained molecular dynamics simulation of proteins. Our approach utilizes a tree-structured representation of collective variables, enabling reconstruction of protein Cartesian coordinates with high fidelity. The evolution of configurations is constructed using a deep neural network trained on trajectories generated from conventional all-atom MD simulations. We demonstrate the framework’s effectiveness using the 168-amino protein target T1027 from CASP14. Statistical distributions of the collective variables and time series of root mean square deviation (RMSD) obtained from our coarse-grained simulations closely resemble those from all-atom MD simulations. This method is not only useful for studying the movements of complex proteins, but also has the potential to be adapted for simulating other biomolecules like DNA, RNA, and even electrolytes in batteries.
@@ I Introduction
Within the bottom-up paradigm of CG methods, researchers leverage the framework of statistical mechanics to define effective interactions, such as potential energiesLiwo et al. (2001). Over the past decades, several traditional coarse-grained methods have emerged, demonstrating significant improvements in capturing protein dynamicsMim et al. (2012); Chu and Voth (2005); Yu et al. (2021); Tóth (2007); Li and Takada (2010). The incorporation of machine learning (ML) techniques has further enhanced the performance of CG simulationsYu et al. (2021); Wang et al. (2019); Husic et al. (2020). For instance, Zhang et al. employed deep neural networks (DNNs) to approximate potential energy surfaces from coarse-grained coordinates, achieving consistent distributions of CVs compared to those obtained from original MD simulationsZhang, Wang, and Weinan (2018). This framework further enabled enhanced sampling of large proteins Wang et al. (2022).
@@
In this research, we present a unified and recursive framework for constructing a bidirectional map between protein Cartesian coordinates and a minimal set of interpretable coarse-grained CVs. This framework incorporates all heavy atoms in the backbone and side chains, enabling a comprehensive representation of the protein structure. Notably, our approach allows MD simulations to be performed solely using the CVs, eliminating the need for explicit Cartesian coordinates. This significantly accelerates simulations while maintaining consistent performance with all-atom MD simulations, as measured by CV distributions and RMSDs at most time steps. Compared to existing methods that achieve only overall statistical consistency on CVsZhang, Wang, and Weinan (2018), our framework offers the advantage extra high accuracy in time series of RMSDs.
@@ II.1 Collective variables
Our protein representation begins with a general description of atomic coordinate hierarchy in chemical compounds. For clarity, this work adopts a non-standard notation. The symbol will represent any dihedral angle and will represent any bond angles. According to Parsons et al.Parsons et al. (2005), the operating matrix for rotating along a normalized rotation vector is
where and . Considering the transformation, the general formula of coordinate transformation goes as a matrix, where is a length-3 vector filled with zeros and is a length-3 vector representing the translation:
Converting the coordinates from the local axes frame (x’y’z’) to parent axes (xyz) can be understood as transforming the parent axes data (axes vector and the origin) to the the local axes data in the parent axes frame, as illustrated in Fig. 1.
@@
where first, translates the frame along x axis by bond length with with being the identity matrix, and ; second, rotates the frame along the new z axis by bond angle with , and ; third, rotates the frame along the new x axis by dihedral angle with , and . These three steps are also illustrated as ①, ② and ③ in the upper pannel of Fig. 1
@@
with index showing the coordinate frames before current frame, and and being the parent and local frame index. is from Eq. 3 in the th coordinate frame. We denote the global operator for converting current coordinates to the "global" reference as
@@
can be readily applied to determine any protein atom’s Cartesian coordinates within the global reference frame, where vector and are original Cartesian coordinates () and extended coordinates (). For convenience, subscript indicates the number and index of coordinate transformation from the "global" coordinate. A superscript of "0" denotes a "local" coordinate within its frame, otherwise it’s a "global" coordinate. As are only determined by the amino type, the Cartesian coordinates are only dependent on . The bond lengths, , used in the calculations can also be retrieved from the standard amino acid structures. It is noteworthy that the the protein backbone C-N bonds have a fixed length of about 1.32 angstroms.
The conversion between a protein’s 3D structure (Cartesian coordinates) and its internal torsion and bond angles is straightforward and is neglected here. This establishes a bidirectional mapping, facilitating analysis from both perspectives. Our calculation only consider the heavy atoms for the sake of computational efficiency. Additionally, rigid moieties like benzene rings require only two angles for complete description. This significantly reduces the number of variables needed, streamlining computational demands. While introducing some additional parameters compared to traditional methods, our approach offers substantial advantages. It allows reconstruction of the complete 3D structure, including side chains, leading to more accurate predictions of interactions – crucial for molecular simulations.
@@ II.1.2 Tree data structure
Recursive data structures naturally lend themselves to representation using tree structures. This approach has been successfully implemented in J. Zhu’s prior work on quantum physics Zhu (2021); Zhu and Scrinzi (2020); Zhu (2020). Similarly, the coordinates of each reference frame within a protein can be effectively represented as a tree structure, as demonstrated in Fig. 1 using an artificial protein with first two amino acids being ILE and TRP as an example. This tree structure efficiently represents protein hierarchy and geometry. Each node serves as a local reference frame, storing the local coordinates of its constituent atoms. Nodes inherit the transformation operator, denoted as (refer to Eq. 5), from their parent nodes (). This inheritance reflects the hierarchical relationships within the protein. Furthermore, the tree captures geometric information. For instance, atoms in TRP forming rigid rings (e.g., CG CH2) reside within a single node. For a more detailed breakdown of the hierarchy within each amino acid, please refer to Table 1 in the appendix. This approach acknowledges their inherent rigidity and reduces the number of redundant parameters needed to define their relative positions, promoting both data storage efficiency and a more natural representation of the protein’s structure.
@@
While the recursive multiplications of matrices can be computationally expensive, the reconstruction of protein 3D structures is only required during trajectory analysis. This means these computationally intensive operations are not necessary for real-time molecular dynamics simulations, thereby minimizing their impact on overall performance. Furthermore, the reconstruction of CV trajectories can be efficiently parallelized across CPUs and GPUs. This distributed processing approach effectively mitigates the performance drawbacks associated with recursive computation time.
II.2 Propagation and molecular dynamics
Simulating large biomolecular systems necessitates capturing their statistical behavior. There exist two main approaches: stochastic methodsHarrison (2009); G Bhanot (1988) and equilibrium samplingKarplus and Petsko (1990); Bussi and Parrinello (2007). Equilibrium sampling uses an initial state that already reflects the desired statistical properties, followed by noise-free simulationsKarplus and Petsko (1990). Stochastic methods introduce random noise at each simulation step to mimic statistical fluctuations, starting from a defined initial state. Examples include random walkCodling, Plank, and Benhamou (2008) and could be decribed by stochastic differential equations (SDE)Slavík (2013). The SDE is also widely applied in computer science for generation of figures Song and Ermon (2019); Song et al. (2021), and later applied molecular dynamics schemes to generate MD trajectoriesWu and Li (2023); Petersen, Roig, and Covino (2021).
This section explores the potential of leveraging strategically chosen data points with minimal correlation from MD trajectories to partially mimic SDEs for statistical analysisCodling, Plank, and Benhamou (2008). By constructing a dataset that reflects the system’s exploration of phase space, we can gain relative information about its dynamics through the lens of an SDE. However, careful selection of the time interval between saved points is crucial, as large gaps can miss important events and hinder accurate analysisSlavík (2013). This approach offers a potentially faster and more efficient way to gain statistical insights into complex systems, particularly when traditional methods become computationally expensive. Our work uses the strategy by utilizing a machine learning propagator trained on saved MD trajectories and this propagator directly predicts the coordinates of next snapshot. We will show that this propagator effectively includes the drift velocity in the SDE.
II.2.1 Propagator network
@@
The network’s task is now to predict the subsequent CVs based on the current ones. These predicted variables are also projected to sine-cosine values, formulated as
@@
where denotes the number of hidden layers in the network, and represent the time indices (timestamps) of the trajectory data. The formula of function in th layer is written as
@@
where and are weights to be learnt, and are activation functions. Our network distinguishes itself from Zhang et al.’s Zhang, Wang, and Weinan (2018) work by introducing a normalization constraint, denoted by , for the sine-cosine operation within the network. This normalization step helps mitigate potential artifacts arising from the projection of angular data into sine and cosine components. While these constraints could also be incorporated into the loss function, we found that applying them directly through significantly accelerates the convergence of the network during training.
@@
The task is to obtain the weights that minimize the loss that can be described as . We use mean squared error (MSE) as the loss function which goes as
@@
where we use error between variables at and those obtained from . This loss can also be extended to a more general form as
where the error computed from subsequent coordinates after are also used. It is easy to see Eq. 13 collapses to Eq. 12 with .
II.2.2 Propagator and SDE
The SDE describes how the CVs evolve over time. Equation
details this evolution, where where represents Gaussian white noise, is the standard deviation dependent on the CVs at time and step size , and denotes the true propagator, which corresponds to the drift velocity in Eq. 18. A subset of loss function in Eq. 13 for predicting the subsequent coordinate based on current time stamp for any time indexes can be written as
where indexes the training data, indicates the number of steps after the current time stamp and subscript refers to the true CVs. The detailed derivation of this loss function can be found in Sec. C.0.1 in the appendix. It can be seen when the loss in Eq. 13 reaches its minimum, the trained DNN effectively approximates the drift velocity represented in Eq. 14. Without the noise term, Eq. 14 collapses to Eq. 10.
Our work builds upon recent advances in SDE-based machine learning, where models learn to reverse a data corruption process (diffusion) Ho, Jain, and Abbeel (2020); Nichol and Dhariwal (2021). However, we propose a novel training scheme due to key differences in our setting. First, data characteristics: Unlike prior works that employ artificially generated, noise-driven trajectoriesHo, Jain, and Abbeel (2020); Nichol and Dhariwal (2021), our model deals with physical, sequential data with inherent physical constraints. Second, Process direction: We learn a forward process directly, contrasting with the reversed process learned in diffusion models. Third, data scarcity: We operate with limited real-world data, as opposed to the large, artificially generated datasets used in previous works. Computing the noises with limited trajectory data may include large numerical errors.
However, we present alternative ways for computing the noises. We make the approximation that at each time stamp of SDE equation, the system stays at equilibrium state and the noise part is formed by the velocities () of the CVs that satisfy the Boltzmann distribution: , where is the equivalent mass, is the Boltzmann constant and is the temperature. Due to the derivation in Sec. C.0.2 in the appendix, all the standard deviation vectors of in the dataset is proportional to a constant vector , which is only dependent on the topology of the protein. Note that the summation of all these standard deviation vectors is proportional to the loss function vector in Eq. 13. Thus we have
where we use individual elements of the vector here compared to summed elements and in Eq. 13. Thus we can use the standard deviation vector in the whole dataset as the noise vector, and select an appropriate coefficient based on the behavior of RMSD in the training data to do the simulation.
II.2.3 Model quality assessment
Due to the inherent uncertainty in the timestep of the SDE, where saved trajectories become uncorrelated after an unknown number of intervals, a range of interval sizes needs to be considered in practical applications. This choice of not only enhances the accuracy of the drift term over a larger timescale but also helps prevent the system from becoming trapped in local minima, when only the propagator operates on the SDE without the noise term for a larger interval . Interestingly, when in Eq. 13, the network demonstrably learns the underlying physical form of the drift force, leading to a more precise simulation, see Sec. D in the appendix. Therefore, we recommend using in practice. We will further substantiate this recommendation with numerical evidence in the application section.
@@ III.1 Protein Structure reconstruction
Fig. 3 depicts the reconstructed initial structure of protein T1027. The reconstructed tertiary structure (cyan) exhibits a near-native match with the original data (blue), highlighting the accuracy of our approach. For comparison, the gray structure was built using all angles except for the bond angles, which were fixed at their theoretical value (). This comparison reveals a significant mismatch, including the absence of part of the alpha helix. This demonstrates the substantial impact of even small variations in bond angles on protein structure, emphasizing the importance of including them in our representation. The secondary structure representation also shows good agreement between the reconstructed structure and the original data, with only minor deviations observed at the ends of a few amino acid side chains. Specifically, analysis of CA atom distances reveals a maximum deviation of 0.26 angstroms, with an average value of 0.04 angstroms. The maximum difference for side-chain atoms is 0.6 angstroms, with a mean of 0.26 angstroms. It is noteworthy that we utilize local coordinates (denoted as - see Section II.1) for each amino acid, retrieved from a separate database independent of the T1027 structure. This approach can introduce slight error accumulation along the protein chain backbone, which can then propagate to the side chains (whose positions are highly dependent on the backbone). However, as will be shown in subsequent sections, the errors introduced by our coarse-grained representation remain consistently below 0.2 angstroms in practical applications, making them effectively negligible.
@@ III.2 MD trajectories prediction
We employed the GROMACSVan Der Spoel et al. (2005); Lindahl, Hess, and van der Spoel (2001); Berendsen, van der Spoel, and van Drunen (1995) software (version 2022.2) for the MD simulations, utilizing a timestep of 1 femtosecond (fs) and saving trajectories every 0.05 nanoseconds (ns). To enhance the efficiency of the simulations, we implemented an in-house developed enhanced sampling method, parallel continuous simulated tempering and a structure based model (PCST+SBM) within GROMACS. Using SBM focuses on the structural variations near a reference structure. This is because comprehensively analyzing all possible protein configurations is computationally demanding for large proteins, and this paper prioritizes a different aspect of the analysis. For a detailed description of PCST+SBM, please refer to references Zhang and Ma (2010); Zang et al. (2014, 2018); Ma et al. (2018). We start the MD simulations from the refinement target in CASP14. Two copies are simulated in PCST, whose temperature distribution is defined by and , respectively. The replica exchange is performed every 5 ns and the time step for updating temperature in Langevin equation is 0.01 ps. In configuring the SBM method, we selected the starting target structure as the reference to promote initial-state fidelity during the MD simulations. This ensures the dynamics primarily focus on exploring configurations near the initial structure. Compared to the original SBM implementation, we introduce a targeted constraint strategy. We constrain only the "slow" residues, which are crucial for calculating the RMSDs. Residues within loops are left unconstrained to facilitate efficient exploration of the conformational space.
@@
This coarse-grained MD simulation utilizes saved trajectories from a prior all-atom MD simulation. The all-atom simulation employed a timestep of 1 fs with data saved every 0.05 ns for a total simulation length of 250 ns. Since this study focuses on replicating the time-series behavior of the simulation, utilizing the full configuration space data becomes unnecessary. While SBM inherently doesn’t provide access to this complete data, obtaining it for such a long protein would also be computationally expensive. Our focus on time-series allows us to achieve a more precise result without requiring the full configuration space. For training the deep neural network, we employ a common data split strategy: 70% of the processed trajectories are used for training, 20% for validation, and the remaining 10% for testing. We configure the network architecture with three hidden layers () and utilize two weight matrices () of size and (details on weight matrix dimensions might be specific to your field).
@@
Our DNN exhibits remarkable efficiency. It can generate a 25ns trajectory using only the initial structure as input and applying the network recursively to predict subsequent coordinates. This process takes merely 0.59 seconds on a standard NVIDIA GeForce GTX 1660 Ti (6G) GPU. In stark contrast, traditional MD simulations on a 256-CPU supercomputer require roughly 20 hours to generate a comparable trajectory. This translates to an impressive speedup of approximately 10,000 times achieved by the DNN network. Even when considering the training time (around 2 hours), the coarse-grained DNN network offers a significant speedup factor of 10 for the entire process. It’s important to note that training can be further optimized by leveraging parallelization techniques and better GPUs.
@@
We first do simulations simply with our DNN propagator using Eq. 10 to show the accuracy of predicted MD trajectories using the test dataset. Fig. 4 depicts the predicted performance of the MD simulation. The upper panel shows the angles (torsion angles and bond angles ). As expected, the angles (represented in the first half) exhibit a wider distribution range due to their inherent flexibility in the protein backbone. In contrast, the angles (, shown in the second half) primarily cluster around 1.23 radians () with a smaller variance ( 0.4 radians) during the simulation. This finding reinforces the earlier observation of bond angle fluctuations within MD simulations. Notably, a small population of angles around radians likely originates from the carbonyl oxygen atoms in the backbone. The comparison of the orange (black) and red (predicted) curves reveals a high degree of consistency between the predicted statistical behavior of the angles and the original values. Similar observations can be made when analyzing the sine and cosine components of these angles, as presented in Fig. E.1 in the appendix.
@@
The RMSD values calculated from the predicted CVs closely match those obtained from the original Cartesian coordinates (Fig. 4, black vs. red lines). Our observations align well with previous structural analysis, where deviations were consistently less than 0.2 angstroms and the average difference using angles () was 0.0691 angstroms with a standard error of 0.0418 angstroms. Additionally, the predicted trajectories closely resemble the original data, particularly between 15-25 ns and 0-5 ns for n=2, as shown by the RMSD profiles. This overall match is also satisfied for n=5 with better agreement during 0-5 and around 12 ns, while worse agreement around 10 and 16 ns. A convergence study (Fig. E.2, Appendix) explored additional n values (1, 3). As expected, n = 1 resulted in significant deviations, indicating insufficient information is learned. Increasing n to 2 eliminated these deviations, suggesting successful information capture. Interestingly, n = 3 led to a slight performance decline, while n = 5 offered minimal improvement while potentially causing slower training and overfitting. Therefore, n = 2 is optimal for convergence. The observed convergence behavior aligns with our prior analysis (Eq. 28), where appears sufficient to capture the essential force field information and generate meaningful data for MD simulations.
We subsequently conducted simulations that included noise terms, calculated from Eq. 16, using training data sets. An analogous propagation with noise was carried out for a duration of 250 ns using Eq. 14. The resulting RMSD trajectories are illustrated in Fig. E.3 in the appendix. These trajectories exhibit significant variance in RMSD values. This variability likely originates from two sources: (1) inaccuracies associated with approximating the noise terms, and (2) the artificial noise scaling factor employed. Despite this variance, the RMSD curves consistently reach values below 3Å, indicating proximity to the native protein configuration, a conformation not explicitly present in the training data. This suggests that the method can be effective in locating the folded state of a protein, but careful calibration of the noise term formulation is crucial.
Our CG method achieves a more precise representation of protein dynamics, as evidenced by the close agreement between our predicted and reference RMSD values, which was never achieved in previous CG methodsWang et al. (2022); Zhang, Wang, and Weinan (2018); Yu et al. (2021); Wang et al. (2019); Husic et al. (2020). The superior precision stems from two key features. First, incorporating both bond and dihedral angles allows for near-native reconstruction of the protein backbone, significantly improving its structural representation. Second, including side-chain information enhances the precision of interactions between amino acids, leading to more realistic protein behavior. Furthermore, the ability to perform bi-directional conversion between CG representation and all-atom structure empower our CG simulations to function as a truly independent tool, extending beyond a complementary role to traditional all-atom simulationsWang et al. (2022); Zhang, Wang, and Weinan (2018).